OpenStudy (anonymous):

Can someone walk me through solving y=3\2x-1

OpenStudy (anonymous):

are you trying to solve for x?

OpenStudy (anonymous):

Well I think so! I mean I am graphing. But the problem is solved for y right

OpenStudy (anonymous):

yes, its already solved for y

OpenStudy (anonymous):

Let me just be honest I get to the problem and I see it and I think okay yeah so what do you want me to do with it?

OpenStudy (anonymous):

If you are graphing it, it is a line with a slope of 3/2 and a y intercept of -1

OpenStudy (anonymous):

I am currently learning to use a table of values intercepts and slopes to graph these

OpenStudy (anonymous):

so then y=mx+b maybe

OpenStudy (anonymous):

thats the equation of a line, so in this case m = 3/2 and b = -1

OpenStudy (anonymous):

So then can you walk me through that maybe. I mean from beginning to end? Or else I won't get it!

OpenStudy (anonymous):

And see I know the formulas but don't really know how to use them

OpenStudy (anonymous):

Since it is a line, if you want to graph it you only need to know two points

OpenStudy (anonymous):

Since your intercept is -1, your first point would be (0, -1)

OpenStudy (anonymous):

to find a second point, plug in any value of x to find the corresponding y value ex. if x = 2 y=(3/2)(2) - 1 so y = 2 and your second point is (2, 2)

OpenStudy (anonymous):

then draw a line through (0, -1) and (2, 2) and thats the graph of your function

OpenStudy (anonymous):

That sort of seems familiar cause my instructor said to kill the fraction use the denominator or somehthing

OpenStudy (anonymous):

yes, since the denominator is 2 if you plug in 2 for x, you get a nice number for y and not a fraction

OpenStudy (anonymous):

Okay and of course you are right. I am working on my labs and I have a test on this tomorrow so do you agree with plugging in a number that per say kills off the fraction and I like nice numbers lol

OpenStudy (anonymous):

Yes, while it doesn't really matter which values of x you plug in it is easier dealing with whole numbers

OpenStudy (anonymous):

Yes it is lol You are a life saver thank you

OpenStudy (anonymous):

OK so what do you say to something that isn't solving for y like y-5x=0

OpenStudy (anonymous):

just solve it for y y = 5x which is a line with slope 5 and intercept 0

OpenStudy (anonymous):

Sorry I don't understand how is that is that according to the slope intercept form?

OpenStudy (anonymous):

y = mx + b so when y = 5x m = 5 and b = 0 so slope 5, intecept 0

OpenStudy (anonymous):

Yeah it worked thanks again

OpenStudy (anonymous):

New Question